### v1

# Data-Driven Loss Functions for Inference-Time Optimization in Text-to-Image Generation

Abstract
Text-to-image diffusion models can generate stunning visuals, yet they often fail at tasks children find trivial—like placing a dog to the right of a teddy bear rather than to the left. When combinations get more unusual—a giraffe above an airplane—these failures become even more pronounced. Existing methods attempt to fix these spatial reasoning failures through model fine-tuning or test-time optimization with handcrafted losses that are suboptimal. Rather than imposing our assumptions about spatial encoding, we propose learning these objectives directly from the model’s internal representations.
We introduce Learn-to-Steer, a novel framework that learns data-driven objectives for test-time optimization rather than handcrafting them. Our key insight is to train a lightweight classifier that decodes spatial relationships from the diffusion model’s cross-attention maps, then deploy this classifier as a learned loss function during inference. Training such classifiers poses a surprising challenge: they can take shortcuts by detecting linguistic traces rather than learning true spatial patterns. We solve this with a dual-inversion strategy that enforces geometric understanding. Our method dramatically improves spatial accuracy: from 0.20 to 0.61 on FLUX.1-dev and from 0.07 to 0.54 on SD2.1 across standard benchmarks. Moreover, our approach generalizes to multiple relations and significantly improves accuracy.
1 Introduction
Figure 1: Learn-to-Steer learns how spatial relationships are encoded in attention maps to guide generation. a) Correctly renders all four orientations (above/below/left/right)  b, c) Handles complex scenes with multiple spatial relationships of up to five objects and three relations. Prompts are for illustration purposes only, see  Sec. 4.3 for actual prompt structure.
Imagine asking a text-to-image model to generate “a dog to the right of a teddy bear.” What could be simpler? Yet surprisingly often, you might get a dog to the left, just a teddy bear with no dog at all, or perhaps a chimeric hybrid of them both. When we venture into more unusual territory like “a cow above a boy” the situation only gets worse. Despite significant progress in text-to-image generation, where models like FLUX.1-dev [26] can create high-quality images from text descriptions, they still struggle with understanding and depicting spatial relationships between objects.
Current approaches to improving text-image alignment generally fall into two camps. Some methods fine-tune the generative model  [17, 50, 6] or train adapter layers [30, 9], risking degradation of the model’s general capabilities. A more flexible paradigm is test-time optimization, which steers the generation process by optimizing the model’s internal latent codes using a loss function. Current methods  [7, 1, 38, 29, 35, 16] typically rely on handcrafted loss functions applied to the model’s attention maps. But here’s the catch: these loss functions rely on simple heuristics that may be suboptimal and address only narrow failure modes, like comparing centers of mass between object attention maps. This raises a fundamental question: instead of imposing our assumptions about how spatial relations should be encoded in attention maps, what if we could learn directly from the model’s own internal representations how to steer it toward better alignment?
Spatial relations represent one type of text-to-image misalignment within a broader family of failures, including incorrect attribute binding, entity neglect, and object counting errors. In this work, we focus specifically on 2D spatial relationships (e.g., ”above”, ”below”, ”left of”, ”right of”) and propose a novel framework, Learn-to-Steer, that learns loss functions for test-time optimization. The core of our method is to treat the steering problem as a learning problem. In an offline first step, we train a lightweight classifier that learns to decode spatial relationships directly from the cross-attention maps produced by the diffusion model. This classifier then serves as a data-driven loss function during inference, providing gradients to steer the latent code toward faithful representation of spatial relations. By learning rather than hand-crafting these guidance signals, we can leverage the full complexity of how diffusion models encode spatial relationships across their internal representations, without needing to manually specify which patterns matter.
Our approach reveals a surprising challenge we term “relation leakage”: when naively trained, classifiers can achieve high accuracy by detecting linguistic traces of relation words in attention maps rather than actual spatial patterns. This creates a unique paradox: we need to predict “left of” while remaining invariant to textual traces of “left of”–essentially learning to extract the visual manifestation of spatial concepts while ignoring their linguistic fingerprints. We address this through a simple yet effective data augmentation strategy that forces the classifier to become invariant to these linguistic artifacts and focus solely on the geometric patterns.
Overall, using learned losses for steering generation significantly improves spatial relation accuracy across multiple diffusion models—from 0.07 to 0.54 on SD2.1 and from 0.20 to 0.61 on FLUX.1-dev – while requiring no model fine-tuning, and avoiding the degradation often seen with fine-tuning approaches. Moreover, to the best of our knowledge, we are the first to demonstrate support for multiple simultaneous spatial relations, directly from input prompts. As shown in Fig. 1, it handles prompts with three relations among five objects. We quantify multiple relations generation accuracy using an extended evaluation scheme.
Our main contributions are: (1) We propose test-time steering using a learned loss function rather than handcrafted losses. (2) We identify and solve the “relation leakage” problem, enabling effective training on cross-attention maps. (3) We demonstrate significant improvements over handcrafted losses, on standard benchmarks (GenEval and T2I-CompBench) across four different diffusion models without any model fine-tuning. (4) We handle multiple simultaneous spatial relations in a single image, demonstrating generalization beyond training on single relations. (5) We introduce an extended evaluation scheme for quantifying multiple relation generation, which we will make publicly available.
2 Related Work
Approaches to improve text-image alignment largely fall into two categories: those that modify the generative model itself, and those that guide it during inference.
The first strategy involves fine-tuning the model’s weights. Some methods fine-tune the generator, for instance by using curriculum-based contrastive learning (EVOGEN [17]), injecting token-ordering information with a curated spatially-aware dataset (COMPASS [50]), re-captioning datasets with spatial phrases (SPRIGHT [6]), or leveraging chain-of-thought with multimodal LLMs (Image-Generation CoT [52]). Others optimize prompt embeddings directly (RRNet [48]). A related approach involves training lighter adapter modules that inject new information, such as bounding boxes (GLIGEN [30]), LLM-generated layouts (Reason-out-Your-Layout [9]), or conditioning maps like Canny edges [5] (ControlNet [51]). In contrast, our work keeps the original pretrained model frozen, preserving its capabilities.
The second major paradigm, and the one our work builds upon, is test-time optimization. Here, the latent code is steered during the denoising process using a guidance loss. However, the dominant approach is to rely on handcrafted loss functions based on heuristics about how attention maps should behave. This includes methods that impose manually designed constraints on self- or cross-attention maps [7, 1, 38, 29, 35, 16], as well as those using explicit layouts as alignment targets [4, 47, 8, 37]. Some methods optimize only the initial noise $z_{T}$ rather than the full trajectory: InitNO [15] uses handcrafted attention-based losses, while ReNO [10] leverages external reward from human preference models. NPNet [53] takes a different approach by training a network to predict initial noises from a dataset of noise-image pairs filtered by human preference models, focusing on overall image quality rather than specific compositional aspects like spatial relations. Our work presents a fundamental departure: instead of hand-engineering objectives or relying on external models, we learn the function to optimize directly from the model’s own internal representations, allowing us to capture the often complex, patterns of spatial encoding.
Other strategies leverage powerful, external VLMs or MLLMs to provide feedback, using their scores to refine or rank generated images [45, 27, 28]. In contrast, we use a small, lightweight classifier trained on the generator’s own internal signals, avoiding dependence on large, black-box models.
A concurrent work, Flow-GRPO [31], fine-tunes the diffusion model directly using the GenEval [14] benchmark as a supervision signal. Since this supervision is not differentiable, they take a reinforcement learning approach using GRPO. Flow-GRPO achieves high quality generations in several categories in GenEval. Naturally, this fine-tuning method involves more complex training and requires substantially more computational resources than our method.
Multiple Spatial Relations.
Generating scenes with multiple spatial relations is a long-standing challenge. Recent benchmarks on simple geometric shapes confirm that even state-of-the-art text-to-image models struggle with this task [44]. A common strategy is to generate an intermediate layout to guide the process. Early works used GNNs on scene graphs to produce the intermediate layout, and were often limited to simple synthetic shapes [18, 3, 23, 11]. More recent layout-guided methods use LLMs to interpret free-form text, yet their focus is typically on object counts and attributes rather than multiple explicit spatial relations [43, 25]. While compositional diffusion approaches can handle multiple relations, they have primarily been demonstrated on simple synthetic shapes [32]. In contrast, our method demonstrates this capability directly from a text prompt on realistic images, without requiring an explicit intermediate layout.
3 Preliminaries: Test-time optimization in text-to-image models
While text-to-image diffusion models produce high-quality images, they often struggle with complex prompts or fail to represent all specified concepts. Test-time optimization addresses these limitations by dynamically steering the generation process toward desired objectives without requiring model retraining. The method works by intervening in the model’s iterative denoising process, which generates an image by progressively removing noise from an initial random latent $z_{T}$ to the final latent representation of the image $z_{0}$.
At each timestep $t$ of this process, the model’s transformer blocks produce a set of cross-attention maps $A_{t}\in\mathbb{R}^{h\times w\times N}$, where $h\times w$ is the spatial resolution and $N$ is the number of text tokens. Each map $A_{t}[:,:,i]$ indicates how token $i$ influences each spatial location, forming a bridge between words and image regions.
Test-time optimization uses these maps to compute a loss $\mathcal{L}(A_{t})$ that quantifies alignment with the desired objective at selected timesteps $t\in\mathcal{T}$, then updates the latent via gradient descent: $z^{\prime}_{t}=z_{t}-\alpha_{t}\nabla_{z_{t}}\mathcal{L}$. While any differentiable signal could be used, cross-attention maps have emerged as the primary optimization target due to their interpretability and direct connection to text-image alignment.
4 Method
Text-to-image diffusion models excel at generating high-quality images, yet they often struggle with accurately representing relationships between objects. While prior work has addressed these alignment challenges through handcrafted loss functions, such approaches are inherently limited: they require domain expertise to design, may encode sub-optimal heuristics, and often target narrow failure modes. What if, instead of manually engineering these losses, we could learn them directly from the model’s own internal representations?
We propose exactly that: a method that learns to interpret and guide diffusion models by understanding how relationships are encoded in their cross-attention maps. Our approach consists of two stages. First, we train a relation classifier that learns to decode spatial relationships from cross-attention patterns. Second, we employ this classifier as a learned loss function during test-time optimization, steering the generation process toward faithful representation of the prompted relationships.
Figure 2: Classifier training pipeline. Given a spatially-aligned image, we perform dual image inversion (positive and negative prompts) to prevent relation leakage. During denoising, we extract the relevant attention maps and use them to train our classifier.
Figure 3: Test-time optimization pipeline. During inference, we extract the relevant cross-attention maps when denoising $z_{t}$ and evaluate their relationship using our trained relation classifier. We then update the latent noise with backpropagation.
4.1 Learning to Decode Relations from Attention
The foundation of our approach is a classifier $\mathcal{C}$, taking cross-attention maps from two objects and predicts their spatial relation. Training such a classifier reveals surprising properties of how diffusion models encode information.
The Relation Leakage Problem.
Consider the seemingly straightforward task of creating training data. Given an image showing “a dog to the left of a cat”, we could invert it using this descriptive prompt, extract the attention maps for “dog” and “cat”, and label them as “left of”. However, our early experiments revealed a critical flaw: classifiers trained this way achieved high accuracy during training but failed when used to guide generation.
Through careful investigation, we discovered the cause: the classifier was learning to take a shortcut. Rather than analyzing the spatial patterns in the attention maps, it was detecting subtle traces of the relation word (“left”) implicitly encoded within them. This phenomenon, which we term relation leakage, meant the classifier could predict “left of” simply by recognizing linguistic artifacts rather than visual evidence.
To verify this we inverted images using deliberately incorrect prompts – for instance, describing a dog that is actually left of a cat as being above it. Remarkably, the classifier predicted “above” rather than the true relation “left of”, confirming that it was reading the prompted relation from the attention maps rather than the actual spatial configuration.
Learning Invariant Spatial Patterns.
To address relation leakage, we need the classifier to learn representations invariant to the traces of the relation word. This presents a unique challenge: unlike typical invariance learning scenarios where we want to ignore a protected attribute while predicting something else  [13, 49, 2], here the “protected attribute” (the relation word traces) and our prediction target (the spatial relation) are semantically the same. We need to predict “left of” while being invariant to linguistic traces of “left of” in the attention maps–essentially learning to extract the visual manifestation of a concept while ignoring its textual fingerprint.
To achieve this invariance, our key insight is to force the classifier to learn from contradictory linguistic signals through a simple, and effective data augmentation strategy. For each training image, we perform inversion twice. First, with a “positive” prompt that correctly describes the spatial relation: “A dog to the left of a cat.” Second, with a “negative” prompt using a randomly sampled incorrect relation: “A dog above a cat.” Crucially, we label both sets of extracted attention maps with the ground-truth relation—“left of”—based on the actual spatial arrangement in the image.
This dual-inversion strategy solves the leakage problem: since the same spatial configuration now appears with different relation words in the prompts, the classifier cannot rely on linguistic shortcuts. It must learn to rely on geometric patterns rather than textual artifacts, thereby achieving invariance to the relation leakage.
The training process is illustrated in Fig. 2, while the full test-time optimization pipeline is shown in Fig. 3. The detailed architecture of our relation classifier is described in Appendix  A.4.
Training Data: Real and Synthetic Images. Our training dataset combines real images from GQA [21] with synthetically generated images from Image-Generation-CoT[52]. Real images provide naturalistic complexity and diverse spatial configurations but yield noisy attention maps when many objects are present. To complement this, we include synthetic images that contain simpler scenes with fewer objects. These images produce cleaner attention maps that resemble and better align with the images encountered during generation.
4.2 Test-Time Optimization with Learned Steering
During generation, the classifier serves as a learned loss function. Given a prompt, we parse a single relation triplet for the subjection, relation and object $(s,r,o)$; and at timestep $t$, aggregate the corresponding attention maps to obtain $\mathbf{A}_{s}$ and $\mathbf{A}_{o}$. The classifier predicts $\hat{p}=\mathcal{C}(\mathbf{A}_{s},\mathbf{A}_{o},t)$, where $\hat{p}$ is the predicted probability distribution over relation classes. We compute a cross-entropy loss against the desired relation $r$ from the prompt:
$\mathcal{L}_{t}=-\log\hat{p}[r],$
and update the latent code with its gradient: $z_{t}\leftarrow z_{t}-\alpha\nabla_{z_{t}}\mathcal{L}_{t}$. We optimize the first 50% of the denoising steps. In FLUX.1-schnell, we only optimize the initial noise.
4.3 Multiple Spatial Relations
A key finding is that our loss, despite being trained only on single pairs of objects, generalizes effectively to prompts that specify multiple spatial relations – ranging from two relations among three objects to more complex cases with 4-5 objects and up to three relations. To accommodate such prompts, we alternate the optimization target at each denoising timestep. For each relation in the prompt, we perform a dedicated optimization step, updating the latent code for one relation before proceeding to the next.
Prompt phrasing has a noticeable impact on performance. While Fig. 1(b) shows structured prompts like “a frog above a sneakers; and the same sneakers below a teapot” for illustrative purposes, we found that simpler, chained prompts of the form “A relation B relation C” are more effective and yield superior results. For example, “a frog above a sneakers below a teapot”.
4.4 Implementation Details
We describe here the main implementation details. Additional details are provided in Appendix A.
Cross-Validation. All architectural choices and hyperparameters were selected through systematic validation on the downstream generation task. We created a validation set using the GenEval framework, carefully ensuring no overlap with test prompts. This validation set contains prompts with similar structure but different object combinations and spatial relations.
Data. For Flux-based models, we use 20K real images from GQA and 4K synthetic images. For each image we extract the attention maps roughly every 5 denoising steps, and including the last (cleanest) step e.g. $\{5,10,15,20,25,49\}$ for multi-step models, and in FLUX.1-schnell we use $\{1,3\}$. For SD-based models, we use a smaller set of training data, because the input dimension is lower ($7\times$ less maps in SD) – with a balanced mixture of 4K real and 4K synthetic images. The GQA dataset undergoes preprocessing to reduce ambiguity: we filter out images with multiple instances of the same object category and remove hypernym relations using WordNet [36]. We use inversion techniques to extract the attention maps of the training data for the denoising steps. For Flux-based models, we use RF-Inversion [40] and for SD-based we use Fixed-Point Inversion [42], using a guidance-scale of 0 for both inversion methods.
Relation Extraction. Relation triplets are extracted from the prompt using SpaCy [19] a natural language parsing library, with some hard-coded exceptions, like ”fire hydrant” or ”hot dog”.
Figure 4: Comparison using FLUX.1-dev (left) and SD 2.1 (right) as base models with prompts from the GenEval [14] benchmark. For each prompt, the same seed is used for all methods.
Figure 5: Common base model failures. Our method addresses common text-to-image generation failures, including incorrect object placement (left), object neglect (middle), and fused, chimeric hybrids (right).
Figure 6: Comparisons with all four base models. Prompts were taken from GenEval [14]. Each pair uses the same seed.
5 Experiments
We compare Learn-to-Steer with strong baselines through quantitative and qualitative evaluation using two benchmark datasets and across four text-to-image diffusion models Flux.1-schnell [26], Flux.1-dev [26], SD2.1 [39] and SD1.4 [39]. Our method demonstrates significant improvements in spatial relations representation than baselines, which often suffer from entity neglect, remaining effective even for atypical scenarios like “A cow above a boy”. Finally, we study the properties of Learn-to-Steer through an ablation study.
5.1 Quantitative Analysis
Method G. Pos. ↑ T. Spatial ↑
FLUX.1-schnell
Vanilla FLUX.1-schnell 0.26 0.199
FOR-Spatial† [24] 0.35 0.238
Learn-to-Steer (Ours) 0.52 0.309
FLUX.1-dev
Vanilla FLUX.1-dev 0.20 0.177
FOR-Spatial† [24] 0.49 0.357
COMPASS [50] 0.60 -
Learn-to-Steer (Ours) 0.61 0.392
SD 2.1
Vanilla SD 2.1 0.07 0.089
FOR-Spatial† [24] 0.27 0.211
FOR [24] 0.34 0.192
STORM† [16] 0.19 0.111
RRNet [48] 0.09 0.106
COMPASS [50] 0.51 -
Learn-to-Steer (Ours) 0.54 0.365
SD 1.4
Vanilla SD 1.4 0.05 0.057
FOR-Spatial† [24] 0.26 0.228
FOR [24] 0.30 0.169
STORM† [16] 0.18 0.101
InitNO [15] 0.12 0.062
COMPASS [50] 0.46 -
Learn-to-Steer (Ours) 0.36 0.278
Table 1: Comparison of spatial relation scores from the GenEval [14] (G. Pos.) and T2I-CompBench [20] (T. Spatial) benchmarks grouped by base diffusion model. All results were computed using the same test set and hardware, except for COMPASS [50], which we report directly from the official paper. †  indicates methods that use handcrafted spatial loss functions.
Baselines: We compare with several methods. (1) COMPASS [50] fine-tunes the model with spatially-aligned data and a text token ordering module. (2) RRNet [48] optimizes prompt embeddings using spatially-aligned reference images. (3) FOR [24] employs a two-stage test-time approach: first optimizing latents $z_{t}$ with handcrafted losses, then refining initial noise $z_{T}$ using VLM feedback. (4) FOR-Spatial [24] uses only the handcrafted spatial loss component from FOR’s first stage for test-time optimization–since FOR only supports SD1.4 and SD2.1, we adapt this component for Flux models by aggregating cross-attention maps per object (details in Sec. A.1). (5) STORM [16] uses handcrafted test-time losses based on optimal transport for spatial relations. (6) InitNO [15] uses handcrafted attention-based loss functions for test-time optimization, focusing on entity neglect and attribute binding failures. We reproduce results for all methods using their shared code, except COMPASS code was not released–we compare with images from their paper in Appendix B.
Benchmarks and Metrics We evaluate spatial alignment using two benchmarks, each with its own automated evaluation protocol. (1) GenEval [14] provides 100 prompts covering four spatial relations: {”above”, ”below”, ”left of”, ”right of”}. We generate four images per prompt and report the percentage of spatially-accurate images among 400 total. Its evaluation metric (G. Pos.) verifies these by comparing the centroids of detected bounding boxes. (2) T2I-CompBench [20] contains prompts for seven spatial relations: {”on side of”, ”next to”, ”near”, ”on the left of”, ”on the right of”, ”on the bottom of”, ”on the top of”}. We extract 216 prompts matching our four trained relations, generate ten images each, and report the percentage of spatially-accurate images. Its metric (T. Spatial) also uses bounding box centers, but applies stricter geometric rules involving the dominant axis of separation and low IoU.
Quantitative Results: Table 1 reveals how learned objectives compare to handcrafted ones across four different models. On SD2.1, spatial accuracy rises from 0.07 to 0.54 on GenEval and from 0.089 to 0.365 on T2I-CompBench—transforming a model that barely works for spatial tasks into one that succeeds more than half the time. FLUX.1-dev shows similar patterns, improving from 0.20 to 0.61 and 0.177 to 0.392 respectively. The results highlight two key findings. First, learned objectives consistently outperform handcrafted spatial losses: STORM achieves 0.19 on SD2.1 GenEval while our method reaches 0.54, and our approach shows similar advantages over FOR-Spatial across all tested models. Second, test-time optimization is competitive with state-of-the-art fine-tuning: COMPASS achieves 0.60 on FLUX.1-dev GenEval through model retraining, while our approach reaches 0.61 by steering the frozen model. Importantly, this pattern holds across architectures—from UNet-based SD models to MMDiT-based Flux variants. Notably, while COMPASS improves spatial relations, it degrades other capabilities—on SD2.1 color accuracy drops from 0.85 to 0.71 and counting accuracy falls from 0.44 to 0.20 (Appendix B). Our approach avoids these trade-offs by steering only when needed.
Multiple Spatial Relations: To evaluate generalization to multiple spatial relations, never seen during training, we generated prompts ranging from two relations among three objects to more complex cases with 4–5 objects and up to three relations. Objects were sampled from a set of 17 common object categories (inspired by  [12] and  [41]). We sampled 300 such multi-relation prompts and generated one image per prompt. Since current benchmarks do not support automatic verification of multiple relations, we extended the GenEval framework to support open-vocabulary evaluation. See implementation details in Sec. A.3. Table 2 evaluates how well FLUX.1-dev and our method handle prompts with multiple simultaneous relations. The base model struggles almost completely, with accuracies near 0.01-0.06 across all settings. In contrast, our method achieves substantial gains: 0.28 on 3 objects with 2 relations, and 0.12 and 0.10 on harder 4- and 5-object cases with 3 relations. These results highlight two findings. First, test-time optimization allows our approach to generalize beyond single-relation training, handling prompts with up to three relations and five objects. Second, these results suggest that relations are handled with some degree of independence, allowing our test-time compositional reasoning to remain robust as complexity increases. While the 0.28 accuracy for two relations is lower than the ideal $(0.61)^{2}=0.372$ for perfectly independent relations, the performance degrades predictably rather than failing completely. For instance, the 0.12 accuracy for three relations is closely approximated by extrapolating from the two-relation result ($0.28\times 0.28/0.61=0.128$). See Sec. A.3.1 for the list of objects and Fig. 9 for additional qualitative comparisons to the base model.
Objects Relations FLUX.1-dev + Ours
3 2 0.02 0.28
4 2 0.06 0.24
4 3 0.01 0.12
5 3 0.01 0.10
Table 2: Automatic evaluation of multiple spatial relations. All results were computed using 300 prompts and the same seed across methods. Bold indicates the best result.
Runtime: We analyze the computational costs of our approach. Our method involves an offline training phase followed by test-time optimization during inference. Classifier training requires 7 hours on an NVIDIA A100 80GB GPU. During inference, test-time optimization introduces computational overhead that varies by model: FLUX.1-schnell increases from 0.5s to 16.5s per image, FLUX.1-dev from 11s to 6 minutes, SD2.1 from 4.5s to 90s, and SD1.4 from 4.5s to 80s. While this overhead is substantial, it enables significant spatial accuracy improvements without requiring model fine-tuning.
5.2 Qualitative Analysis
Figures 4 and 6 show qualitative examples across four base models and several baselines. Compared to other methods, our approach consistently generates high-quality images that accurately reflect the spatial relations in the prompts. It also handles generating atypical and uncommon scenarios well, as can be seen in  Fig. 4 with “A bus below a toothbrush” and in  Fig. 6 with “An elephant below a surfboard. In contrast, other baselines tend to neglect entities: in Fig. 4, none depict the vase from “A vase above a fire hydrant”, and the zebra or the chair are missing from “A chair to the left of a zebra”. These examples, along with Figure 5, demonstrate how our method overcomes the three main failure modes: incorrect object placement, entity neglect, and object fusion.
In Appendix B, qualitative comparisons with COMPASS show we achieve equivalent spatial accuracy while maintaining style and aesthetics closer to the base model. For additional qualitative results, see Appendix  C, which includes extensive comparisons on both GenEval (Figures 11, 12, 13, and 14) and T2I-CompBench (Figures 15, 16, 17, and 18) benchmarks, as well as uncurated generation results in Fig. 10.
5.3 Ablation Study
Component Variant G. Pos. ↑
Learn-to-Steer (Ours) 0.57
Leakage Handling w/o dual inversion (pos only) 0.425
Training Data Real images only (GQA) 0.365
Synthetic images only 0.373
Table 3: Ablation results on FLUX.1-schnell validation set.
We examine the key design choices of our method through systematic ablation on FLUX.1-schnell. Results are summarized in Table 3.
Leakage Handling. Our dual-inversion approach is crucial for addressing relation leakage. Without it, training on positive prompts only yields 0.42.5. The dual-inversion strategy eliminates this leakage, enabling our method to reach 0.57.
Training Data Composition. Real images from GQA alone achieve 0.365, while synthetic images alone reach 0.373. Real images provide naturalistic complexity but yield noisy attention maps in cluttered scenes. Synthetic images produce cleaner attention patterns that better match generation scenarios. Our full method combines both data sources, leveraging their complementary strengths.
6 Limitations
Figure 7: Method Limitations. (a) 2D spatial relations may not align with 3D scene geometry—the spoon appears below the bear in 2D but is in front of it in 3D. (b) Multiple spatial relationships with three relations: When prompted with four objects and three spatial relations, our method is sometimes able to generate matching images. However, typical failure cases—such as entity neglect—still occur
Our method has several limitations. First, we train on 2D spatial relations, which can conflict with 3D scene understanding. As shown in Fig. 7(a), a spoon positioned in front of a bear in 3D space may appear ”below” it in the 2D projection.
Second, handling scenarios with multiple spatial relationships becomes more challenging when three relations are involved compared to two. Our accuracy drops from 0.28 with two relations to just 0.12 with three, showing that as the number of relations increases, accuracy tends to decrease.  Fig. 7(b) shows both successful generations and typical failure cases, such as entity neglect.
Finally, despite substantial improvements, there is still a lot of headroom for improvement. Entity neglect and incorrect positioning persist, as shown in uncurated generation results in Fig. 10 in the Appendix.
7 Conclusions
We presented Learn-to-Steer, a data-driven alternative to handcrafted spatial losses for test-time steering of text-to-image diffusion models. By training a lightweight classifier to decode spatial relationships directly from cross-attention maps and deploying it as a learned loss function, we achieve substantial improvements in spatial accuracy across multiple models without any fine-tuning. A key advantage of our approach is that it preserves the model’s full capabilities by applying steering only when spatial relations are needed, avoiding the performance degradation often seen with fine-tuning methods. While test-time optimization does increase inference time, this selective application ensures that the model’s general generative abilities remain intact for prompts without spatial constraints. This cost enables applications requiring high spatial accuracy—including generating multiple relations data. Our work demonstrates that learning from the model’s own internal representations can be more effective than handcrafted objectives, opening new directions for improving text-to-image alignment.
Acknowledgments
This study was funded by the Israeli Ministry of Science, Israel-Singapore binational grant, and by a grant from the National Committee of Budgeting (VATAT). We thank Yoad Tewel, Ohad Rahamim, and Ben Fishman for providing feedback on an earlier version of this manuscript.
Supplementary Material
A Additional Implementation Details
A.1 Adapting FOR’s handcrafted spatial loss to Flux-based models
We adapted FOR’s [24] handcrafted spatial loss as follows: For each pair of objects, we first aggregated all cross-attention maps across all layers and heads into a single attention map per object. Subsequently, we applied the manual spatial loss function on the two aggregated maps in exactly the same manner as in the official implementation. To ensure a fair comparison, we ran a hyperparameter search over the scale factor of the step size used during test-time optimization to find the optimal setup.
A.2 Benchmarks
To ensure a fair comparison, we normalize all relation names to {”above”, ”below”, ”to the left of”, ”to the right of”} across benchmarks. All baselines we reproduced using this normalized set. We will release normalized test sets for both benchmarks.
A.3 Multiple Relations Evaluation
Objects Relations Manual Eval. Automatic Eval.
3 2 0.29 0.28
4 3 0.12 0.12
5 3 0.07 0.10
Table 4: Comparison of automatic evaluation vs. manual evaluation of multiple spatial relations.
We investigated why current benchmarks do not support automatic verification of multiple relations. Specifically, we focused on the GenEval [14] framework and identified the main culprit as the object detector, which we replaced with CLIPSeg [34]. To enable evaluation, we generated metadata files in GenEval’s expected structure, where each prompt is paired with a list of required objects along with their pairwise relations. To verify that our automatic evaluation is reliable, we compared its results against manual evaluation conducted by one of the authors (Table 4). The scores align closely across all settings.
A.3.1 Evaluation Objects
For the multi-relation evaluation, we used the following 17 object categories (inspired by Textual-Inversion [12] and DreamBooth [41]): backpack, bowl, cat, sneakers, duck, robot, teapot, vase, corgi, mug, furby, teddy bear, frog, mouse, table, chair, and umbrella.
A.4 Architecture Details
First we process each set of attention maps individually, using a transformer based aggregation module that maps each set of maps $L\times H\times h\times w$ to one single $h\times w$ map per image, where $L$, $H$ are the number of cross attention layers and heads and $h$, $w$ are the attention map dimensions. Then we process these two maps jointly to predict the relation class.
Aggregation Module. For a given object, the aggregator receives all $L\times H$ attention maps of size $h\times w$. Each map is treated as a separate token in a sequence of length $L\cdot H$. We then project each token from $h\cdot w$ dimensions to a shared embedding dimension $d=256$. This yields a sequence length $L\cdot H$ with a feature dimension of $d$, to which we add standard positional encodings [46] in $\mathbb{R}^{d}$ that signifies the spatial location. A timestep embedding $\phi(t)\in\mathbb{R}^{d}$ is also added to all tokens, in the input to every transformer layer, allowing the aggregator to adapt to the current denoising stage. A four-layer, four-head transformer then processes this sequence to exchange information across different layers and heads. To aggregate the transformer output from $\mathbb{R}^{L\cdot H,d}$ to $\mathbb{R}^{d}$, we use another four-layer, single-head transformer. For the second transformer, we learn a query vector in $\mathbb{R}^{1\times d}$, while the key and value are projected from the output of the first transformer ($\mathbb{R}^{L\cdot H,d}$). Finally, a linear layer maps the output in $\mathbb{R}^{d}$ to $\mathbb{R}^{h\cdot w}$ reshaped to an aggregated map $\mathbf{A}_{\text{agg}}\in\mathbb{R}^{h\times w}$. The same aggregator (shared weights) processes both subject and object maps.
In practice, to process SD-based maps we use two-layer, two-head transformers instead of four-layer, four-head.
Classification Module. The two aggregated maps are stacked into a two-channel image, flattened, projected to $d=256$ dimensions, tagged with standard positional encodings and passed through a four-layer eight-head Transformer. Finally, the transformer’s output is concatenated with a low-dimension timestep embedding (16-d) and fed to an MLP that returns logits over the relation classes. The complete architecture is trained jointly end-to-end.
In practice, to process SD-based maps we use two-layer, four-head transformers in the classification module.
A.5 Other details
Classifier Training. For training the classifier we used a batch size of 64, with AdamW optimizer [33], a weight decay of $0.05$, and a learning rate of $5\cdot 10^{-5}$ for Flux-based models and $0.0001$ for SD-based models. We used a ReduceOnPlateu scheduler with a factor of $0.5$ and patience $5$. We trained the classifier for a maximum of $200$ epochs in Flux-based models and $150$ for SD-based models, with early stopping patience of $10$ epochs.
Test-Time Optimization and Inference. We optimize the first $50\%$ of the denoising steps. In FLUX.1-schnell, we only optimize the initial noise. The step size $\alpha$ is $5$ for SD2.1 and Flux-based models, and $7.5$ for SD1.4 and number of optimization iterations per step are $15$.
We use the guidance scales suggested by HuggingFace [22], specifically $3.5$ for FLUX.1-dev, 0 for FLUX.1-schnell, and $7.5$ for SD based models.
Image Resolution. We generate images at different resolutions depending on the base model: $256\times 256$ for FLUX.1-schnell and $512\times 512$ for FLUX.1-dev, SD2.1, and SD1.4. Neutral Relation Class. We included a fifth, ”neutral” class during training. This class captures non-directional relationships (described by ”is”) and acts as a background category. We do not use this neutral class during generation.
B Comparison with COMPASS
While fine-tuning methods like COMPASS [50] achieve strong performance on spatial relations, they often come at the cost of degrading other alignment capabilities. Table 5 shows this trade-off using results reported in the COMPASS paper.
Model Position ↑ Count ↑ Color ↑
Vanilla SD2.1 0.07 0.44 0.85
COMPASS 0.51 0.20 0.71
Learn-to-Steer (Ours) 0.54 0.44* 0.85*
Table 5: GenEval results comparing fine-tuning (COMPASS) vs. test-time optimization (Ours) on SD2.1. COMPASS results are from their paper. *Our method preserves the vanilla model’s performance on non-spatial tasks since we only apply steering when spatial relations are present in the prompt.
As shown in Table 5, COMPASS’s spatial relation improvements come with measurable degradation in other capabilities: color accuracy drops from 0.85 to 0.71 and counting accuracy drops from 0.44 to 0.20. In contrast, our test-time optimization approach preserves the original model’s capabilities across all metrics. Additionally, the qualitative comparison in Fig. 8 shows that our method achieves equivalent spatial accuracy while maintaining style and aesthetics closer to the base model.
Figure 8: Comparing COMPASS [50] to our method, using FLUX.1-dev [26] as the base model.
C Additional Qualitative Results
We perform extensive comparisons on both GenEval (Figures 11, 12, 13, and 14) and T2I-CompBench (Figures 15, 16, 17, and 18).
Figure 9: Handling Multiple Relations. As opposed to the base model, our method is capable of generating complex scene structures that contain multiple relations in a single prompt.
Figure 10: Uncurated image generation results from FLUX.1-dev [26] with Learn-to-Steer, using validation prompts from GenEval [14]. Green boxes mark spatially-aligned images (per GenEval), while red boxes mark misaligned ones.
Figure 11: Qualitative comparison using FLUX.1-schnell [26] with prompts from the GenEval [14] benchmark. For each prompt, the same seed is used for all methods.
Figure 12: FLUX.1-dev on the GenEval benchmark
Figure 13: SD 2.1 on the GenEval benchmark
Figure 14: SD 1.4 on the GenEval benchmark
Figure 15: FLUX.1-dev on the T2I-CompBech benchmark
Figure 16: FLUX.1-schnell on the T2I-CompBech benchmark
Figure 17: SD 2.1 on the T2I-CompBech benchmark
Figure 18: SD 1.4 on the T2I-CompBech benchmark